A step ladder has legs of length 2.2m and the maximum distance between the legs is 0.9m. Find the largest possible angle in which the ladder can be opened.
Answer by Alan3354(69443)

A step ladder has legs of length 2.2m and the maximum distance between the legs is 0.9m. Find the largest possible angle in which the ladder can be opened.
------------
If you view it as a triangle, it's isosceles.
1/2 the angle at the top = asin(0.45/2.2) =~ 11.8 degs
--> 23.6 degs
-----------------------
Ladders don't have a sharp points at the top, tho.